# Weak Cosmic Censorship with spinning particles in Kerr-(A)dS spacetimes

### v1

@@ Abstract
We investigate the weak cosmic censorship conjecture by analyzing the dynamics of spinning timelike particles dropped along the rotational axis of an extremal Kerr–(anti-)de Sitter black hole. This idea was first considered in a seminal paper by Wald and later by Needham but both analyses were restricted to asymptotically flat spacetimes. We generalize these studies, involving spinning particles, to rotating spacetimes with non vanishing cosmological constant. We examine whether the absorption of such particles can overspin the black hole beyond extremality, potentially leading to the formation of a naked singularity. In asymptotically de Sitter spacetime, we find that particles that are captured cannot overspin the black hole. For generic initial conditions, similar conclusions hold also with anti-de Sitter asymptotics. However, in this case, we identify a narrow set of finely tuned trajectories, corresponding to particles injected very close to the horizon, that lead to the disruption of the event horizon.
Introduction. The possible appearance of spacetime singularities in General Relativity – regions where Einstein’s field equations break down– poses an important problem in the deterministic evolution of the spacetime geometry [1]. To preserve the theory’s predictive power, it is crucial that such singularities do not influence regions accessible to distant observers. This is the core idea behind the weak cosmic censorship conjecture (WCCC), formulated by Penrose [2]. It posits that, under generic initial conditions and for physically reasonable matter, singularities arising from gravitational collapse are always hidden from view to distant observers by an event horizon.
@@
Realistic BHs, however, are not isolated objects in asymptotically flat spacetime. Observations indicate that our universe is undergoing accelerated expansion, consistent with the presence of a positive cosmological constant and a de Sitter (dS) background [5, 6, 7]. Meanwhile, asymptotically anti-de Sitter (AdS) spacetimes –characterized by a negative cosmological constant– have become central to theoretical physics due to their role in the gauge/gravity duality [8, 9]. Thus, understanding the fate of cosmic censorship in rotating BHs within both dS and AdS backgrounds is of paramount importance. Various studies have examined the conjecture in AdS settings under diverse conditions [10, 11, 12, 13, 14, 15, 16, 17, 18, 19], as well as in the Kerr-dS case [11, 20, 21, 22], generally finding that the WCCC remains intact  [^1] Nevertheless, exceptions have been reported: in particular, charged models in AdS4 have suggested that regions with arbitrarily large curvatures become visible to boundary observers in certain fine-tuned configurations [53, 44]. Additionally, small Kerr-AdS suffer from the superradiant instability, which also might impact the horizon stability and, thus, the WCCC [54, 55, 56, 57]..
In this letter, we generalize Wald’s original thought experiment to assess whether spinning particles, dropped into an extremal Kerr-(A)dS BH along its axis of rotation, can breach the extremality condition, effectively destroying the horizon and violating the WCCC. See Fig. 1 for a pictorial representation of the setup considered.
Despite the extensive exploration of the WCCC via test particles in geodesic motion, studies involving spinning particles are scarce, and have only been carried out in asymptotically flat spacetimes [3, 24], to the best of our knowledge. Additionally, the Sorce-Wald method has not yet been applied to rotating and non-asymptotically flat BHs. Only a handful of results exist in static charged models with a cosmological constant [25, 26, 27].
@@
Furthermore, we restrict the parameters to $10\leq a/L<$ 1, in order to avoid divergences in the line element (1)  [^2] The restriction to non-negative values for $a/L$ is only a matter of convention. Negative values can be considered by flipping the sign of $a$..
In the limit $L^{2}\to\infty$, both Eqs. (6) and (7) reduce to $r_{\text{H}}=M=a$, as expected, since the latter is the extremality condition for Kerr BHs. Spinning test particles along the axis of the Kerr-(A)dS metric. The equations describing the motion of spinning test particles are known as the Mathisson-Papapetrou-Dixon (MPD) equations [34, 35, 36]:
@@
Figure 3: Comparison, for the Kerr-AdS case, between the upper bound on $\hat{s}/L$ (light blue solid curve) allowing the particle to be captured by the BH, given by Eq. (15), the lower bound on $\hat{s}/L$ (violet solid curve) required to destroy the event horizon, as indicated by Eq. (Weak Cosmic Censorship with spinning particles in Kerr-(A)dS spacetimes), and the radius of the extremal BH (7) (black dashed curve), which determines the regime of validity of the point-particle approximation. $a/L$ is restricted to the range $[0,1)$ in accordance with the discussion in the main text. Note that in this case there is an overlap region where the particle can have sufficiently small spin to be captured by the BH but still have sufficiently large spin to destroy the event horizon.
@@
Conclusions and outlook. Our analysis extends the existing literature on testing the WCCC with spinning particles in asymptotically flat spacetimes by exploring analogous processes in the presence of a timelike boundary (AdS) and a cosmological horizon (dS). We find that an extremal Kerr–dS black hole resists overspinning attempts via the absorption of spinning particles falling along its axis of rotation. A similar conclusion holds generically in the Kerr-AdS case. However, we identify finely tuned scenarios in which a particle originating near the horizon leads to its destruction. These findings complement prior results that suggest potential WCCC violations in AdS spacetimes [43, 44], and may offer further insights in holographic contexts [45, 46].
These results also connect with broader considerations of geometric inequalities involving BH parameters. In asymptotically flat spacetimes, the WCCC is closely tied to the Penrose inequality, which relates a BH’s mass and horizon area [47]. A corresponding inequality has been conjectured for AdS spacetimes [48, 49] or in dS (see, e.g., Ref. [50]).
@@
A recent proposal [51] presents a new class of inequalities for stationary asymptotically AdS BHs, relating their horizon area to other thermodynamic quantities. Our results, which highlight the fragility of the Kerr-AdS BH in finely tuned parameter regimes, provide a framework to similarly test the WCCC using more general rotating BHs in AdS, including charges and/or matter fields, and offer a complementary perspective to the limits imposed by the inequalities put forward in Ref. [51]
@@
Finally, our analysis lays the foundation for extending these ideas to rotating and accelerated black holes, which, in combination with the braneworld scenario, admit a holographic dual description in terms of quantum-corrected black holes [52]. In this framework, the injection of a spinning particle in the bulk acquires a lower-dimensional interpretation on the brane, directly connecting with the results of Ref. [19].
@@
Acknowledgments. AMF acknowledges the support from EU Horizon 2020 Research and Innovation Programme under the Marie Skłodowska-Curie Grant Agreement no. 101149470. JVR acknowledges the support from Fundação para a Ciência e a Tecnologia under project 2024.04456.CERN. APS gratefully acknowledges support from a research grant funded under the INFN–ASPAL agreement as part of the Einstein Telescope training program. APS is also partially supported by the MUR FIS2 Advanced Grant ET-NOW (CUP: B53C25001080001) and by the INFN TEONGRAV initiative.
Appendix A Supplementary Material
@@
In the following, we provide additional details of the calculations supporting the main text’s conclusion that spinning particles moving along the axis of rotation of an extremal Kerr-AdS BH can, in certain fine-tuned situations, lead to horizon disruption.
@@
Since for a spinning particle moving along $\theta=0$, the 4-momentum is parallel to $v^{\mu}$ (see the main text), we have also $p^{r}=0$. The timelike constraint $p_{\mu}p^{\mu}=-m^{2}$, thus, reduces to $g_{tt}\left(p^{t}\right)^{2}=-m^{2}$. Evaluating the latter onto the axis of rotation and at $r=r_{\text{inj}}$, considering future directed particles with $p_{t}<0$ and using Eq. (13) in the main text gives
@@
Here, it is assumed that $M$ takes its extremal value (Eq. (7) in the main text). Since the AdS length $L$ can be scaled away and also the particle mass $m$ enters only through the ratios $E/m$ and $s/m$, we can compute the energy-to-mass ratio of the particle depending on $r_{\text{inj}}$, $s/m$ and $a$. The spin-to-mass ratio $s/m$ is subjected to the point-particle constraint, i.e., $s/m\ll r_{\text{H}}$, while we restrict also to $10\leq a/L<$ 1.
@@
and use the fact that $E$ and $m$ are constants of motion. Then, when the above quantity is smaller than $1$, we have automatically WCCC preservation (according to Fig. 3 in the main text). If, instead, there are some particles that satisfy $\hat{s}/r_{\text{H}}>1$, they live above the dashed line in Fig. 3 and potentially can overspin the BH if they reach the overlapping region when $\hat{s}/r_{\text{H}}>\hat{s}^{\text{AdS}}_{2}/r_{\text{H}}$. A necessary condition for this to happen is that also $\hat{s}/r_{\text{H}}<\hat{s}^{\text{AdS}}_{1}/r_{\text{H}}$ must hold, in order to ensure particle absorption by the BH.
@@
We performed a numerical scan of the parameter space to assess whether such regions, where $\hat{s}/r_{\text{H}}>\hat{s}^{\text{AdS}}_{2}/r_{\text{H}}$, exist. We will consider particles with $s/m\leq 0.2\,r_{\text{H}}$ (so that the point-particle approximation, $s/m\ll r_{\text{H}}$, is satisfied). Fig. 4 reports, for different values of $a/L$, the regions in the parameter space where BH horizon disruption occurs. We have checked that $\hat{s}/r_{\text{H}}<\hat{s}^{\text{AdS}}_{1}/r_{\text{H}}$ is always true in the parameter space considered, so all the particles in the considered parameter regions will be absorbed.
@@
Figure 4: Regions in the parameter space $[s/m,\,(r_{\text{inj}}-r_{\text{H}})/r_{\text{H}}]$ (the spin-to-mass ratio and the normalized initial distance of the particle from the horizon, respectively), where the particle has $\hat{s}/r_{\text{H}}>\hat{s}^{\text{AdS}}_{2}/r_{\text{H}}$ and thus can be absorbed and spin the BH past extremality. For each plot, we fixed the value of $a/L$ (determining also the value of $r_{\text{H}}$ through Eq. (7) in the main text) and we restricted to particles with $s/m\leq 0.2\,r_{\text{H}}$, starting (with zero velocity) from a radial distance from the horizon equal to $r_{i}-r_{\text{H}}\geq 10^{-3}\,r_{\text{H}}$. For $s/m=0$, one always has $\hat{s}/r_{\text{H}}<\hat{s}^{\text{AdS}}_{2}/r_{\text{H}}$, consistent with no horizon disruption.
As an example, we consider the numerical parameters of the particle mentioned in the main text: $s/m=0.2\,r_{\text{H}}$, injected at the radius $r_{\text{inj}}=1.001\,r_{\text{H}}$, very close to the BH event horizon with $\tilde{a}=0.1$.
Fig. 5, instead, illustrates the effective potential,
$V_{\text{eff}}=1+\frac{r^{2}}{L^{2}}-\frac{2Mr}{r^{2}+a^{2}}-\left[1-a\hat{s}\left(-\frac{1}{L}+\frac{2Mr}{(r^{2}+a^{2})^{2}}\right)\right]^{2}\,,$ (26)
which governs the radial motion of the particle according to $\dot{r}^{2}+V_{\text{eff}}=0$. Here, the overdot denotes a derivative with respect to the particle’s proper time. The potential is negative in the interval $[r_{\text{H}},\,r_{\text{inj}}]$, demonstrating consistency with particle capture by the BH.
Figure 5: Effective potential (26) for selected values of the particle’s and the BH’s parameters: $s/m=0.2\,r_{\text{H}}$, $r_{\text{inj}}=1.001\,r_{\text{H}}$, $\tilde{a}=0.1$. The horizontal axis starts from the corresponding position of the event horizon. The inset shows a zoom of the near-horizon region. The dashed vertical line corresponds to the turning point $r_{\text{inj}}=1.001\,r_{\text{H}}$ at which the particle is injected.

### v2

@@ Abstract
We investigate the weak cosmic censorship conjecture by analyzing the dynamics of spinning timelike particles dropped along the rotational axis of an extremal Kerr–(anti-)de Sitter black hole. This idea was first considered in a seminal paper by Wald and later by Needham but both analyses were restricted to asymptotically flat spacetimes. We generalize these studies, involving spinning particles, to rotating spacetimes with non vanishing cosmological constant. We examine whether the absorption of such particles can overspin the black hole beyond extremality, potentially leading to the formation of a naked singularity. In asymptotically de Sitter spacetime, we find that particles that are captured cannot overspin the black hole. Similar conclusions hold also with anti-de Sitter asymptotics, but the analysis is more subtle, requiring careful consideration of the point particle approximation.
Introduction. The possible appearance of spacetime singularities in General Relativity, regions where Einstein’s field equations break down, poses an important problem in the deterministic evolution of the spacetime geometry [1]. To preserve the theory’s predictive power, it is crucial that such singularities do not influence regions accessible to distant observers. This is the core idea behind the weak cosmic censorship conjecture (WCCC), formulated by Penrose [2]. It posits that, under generic initial conditions and for physically reasonable matter, singularities arising from gravitational collapse are always hidden from view to distant observers by an event horizon.
@@
Realistic BHs, however, are not isolated objects in asymptotically flat spacetime. Observations indicate that our universe is undergoing accelerated expansion, consistent with the presence of a positive cosmological constant and a de Sitter (dS) background [5, 6, 7]. Meanwhile, asymptotically anti-de Sitter (AdS) spacetimes —characterized by a negative cosmological constant— have become central to theoretical physics due to their role in the gauge/gravity duality [8, 9]. Thus, understanding the fate of cosmic censorship in rotating BHs within both dS and AdS backgrounds is of paramount importance. Various studies have examined the conjecture in AdS settings under diverse conditions [10, 11, 12, 13, 14, 15, 16, 17, 18, 19, 20], as well as in the Kerr-dS case [11, 21, 22, 23], generally finding that the WCCC remains intact  [^1] Nevertheless, exceptions have been reported: in particular, charged models in AdS4 have suggested that regions with arbitrarily large curvatures become visible to boundary observers in certain fine-tuned configurations [49, 48]. Additionally, small Kerr-AdS suffer from the superradiant instability, which also might impact the horizon stability and, thus, the WCCC [59, 60, 61, 62]..
In this letter, we generalize Wald’s original thought experiment to assess whether spinning particles, dropped into an extremal Kerr-(A)dS BH along its axis of rotation, can breach the extremality condition, effectively destroying the horizon and violating the WCCC. See Fig. 1 for a pictorial representation of the setup considered. Refs. [15, 22] proved more generally that extremal Kerr-Newman-(A)dS cannot be overcharged/overspun with test fields obeying the null energy condition. Since physical spinning bodies —of which spinning particles can be considered a limiting case— should satisfy the dominant energy condition [25], and that implies the null energy condition, the analysis we present falls under the conditions of the theorems proved in [15, 22]. Nevertheless, the specific calculation we perform is useful to compare with similar tests in rotating BH backgrounds different from Kerr-(A)dS.
Despite the extensive exploration of the WCCC via test particles in geodesic motion, studies involving spinning particles are scarce, and have only been carried out in asymptotically flat spacetimes [3, 26, 27], to the best of our knowledge. Additionally, the Sorce-Wald method has not yet been applied to rotating and non-asymptotically flat BHs. Only a handful of results exist in static charged models with a cosmological constant [28, 29, 30].
@@
Furthermore, we restrict the parameters to $0\leq a/L<1$, in order to avoid divergences in the line element (1)  [^2] The restriction to non-negative values for $a/L$ is only a matter of convention. Negative values can be considered by flipping the sign of $a$..
In the limit $L^{2}\to\infty$, both Eqs. (6) and (7) reduce to $r_{\text{H}}=M=a$, as expected, since the latter is the extremality condition for Kerr BHs.
Spinning test particles along the axis of rotation. The equations describing the motion of spinning test particles are known as the Mathisson-Papapetrou-Dixon (MPD) equations [37, 38, 39]:
@@
A natural question then arises regarding the fate of particles injected in the immediate vicinity of the BH. In this regime, the finite size of the particle should also be compared with the distance from the horizon. Specifically, it should satisfy $r_{0}<r_{\text{inj}}-r_{\text{H}}$. We find that particles with fine-tuned parameters capable of disrupting the horizon possess a size exceeding this bound, implying they were already inside the BH when they were injected  [^4] We thank J. Natário for pointing out this consideration.. Conversely, particles smaller than $r_{\text{inj}}-r_{\text{H}}$ are unable to disrupt the horizon, no matter how close from it they are injected (see the Appendix for further details).
@@
Figure 3: Comparison, for the Kerr-AdS case, between the upper bound on $\hat{s}/L$ (light blue solid curve) allowing the particle to be captured by the BH, given by Eq. (15), the lower bound on $\hat{s}/L$ (violet solid curve) required to destroy the event horizon, as indicated by Eq. (Weak Cosmic Censorship with spinning particles in Kerr-(A)dS spacetimes), and the radius of the extremal BH (7) (black dashed curve), which determines the regime of validity of the point-particle approximation. $a/L$ is restricted to the range [0,1)[0,1)[ 0 , 1 ) in accordance with the discussion in the main text. Note that in this case there is an overlap region where the particle can have sufficiently small spin to be captured by the BH but still have sufficiently large spin to destroy the event horizon.
@@
Conclusions and outlook. Our analysis extends the existing literature on testing the WCCC with spinning particles in asymptotically flat spacetimes by exploring analogous processes in the presence of a timelike boundary (AdS) and a cosmological horizon (dS). We find that an extremal Kerr–dS BH resists overspinning attempts via the absorption of spinning particles falling along its axis of rotation. Similar conclusions hold in the Kerr-AdS case, although neglecting a careful evaluation of the point particle approximation would lead to the destruction of the horizon in finely tuned scenarios. These results are consistent with the findings of Ref. [15]. Considering different matter content and BH backgrounds, potential WCCC violations in AdS spacetimes have been suggested [47, 48, 49]. At any rate, our results for Kerr-AdS may offer further insights in holographic contexts [50, 51, 52].
@@
These results also connect with broader considerations of geometric inequalities involving BH parameters. In asymptotically flat spacetimes, the WCCC is closely tied to the Penrose inequality, which relates a BH’s mass and horizon area [53]. A corresponding inequality has been conjectured for AdS spacetimes [54, 55] or in dS (see, e.g., Ref. [56]). A recent proposal [57] presents a new class of inequalities for stationary asymptotically AdS BHs, relating their horizon area to other thermodynamic quantities. Our results, which highlight the importance of the point-particle approximation in the Kerr-AdS background, provide a framework to similarly test the WCCC using more general rotating BHs in AdS, including charges and/or matter fields, and offer a complementary perspective to the limits imposed by the inequalities put forward in Ref. [57].
@@
Finally, our analysis lays the foundation for extending these ideas to rotating and accelerated BHs, which, in combination with the braneworld scenario, admit a holographic dual description in terms of quantum-corrected BHs [58]. In this framework, the injection of a spinning particle in the bulk acquires a lower-dimensional interpretation on the brane, directly connecting with the results of Ref. [20].
@@
Acknowledgments. We thank J. Natário for insightful and useful discussions. AMF acknowledges the support from EU Horizon 2020 Research and Innovation Programme under the Marie Skłodowska-Curie Grant Agreement no. 101149470. JVR acknowledges the support from Fundação para a Ciência e a Tecnologia under project 2024.04456.CERN. APS gratefully acknowledges support from a research grant funded under the INFN–ASPAL agreement as part of the Einstein Telescope training program. APS is also partially supported by the MUR FIS2 Advanced Grant ET-NOW (CUP: B53C25001080001) and by the INFN TEONGRAV initiative.
Appendix A Appendix
@@
In the following, we provide additional details concerning the calculations supporting the main text’s conclusion that spinning particles moving along the axis of rotation of an extremal Kerr-AdS BH cannot overspin it, in the regime where the point particle approximation is valid.
@@
Since for a spinning particle moving along $\theta=0$, the 4-momentum is parallel to $v^{\mu}$ (see the main text), we have also $p^{r}=0$ at the turning point. The timelike constraint $p_{\mu}p^{\mu}=-m^{2}$ in this situation, thus, reduces to $g_{tt}\left(p^{t}\right)^{2}=-m^{2}$. Evaluating the latter onto the axis of rotation and at $r=r_{\text{inj}}$, considering future directed particles with $p_{t}<0$ and using Eq. (13) in the main text gives
@@
Here, it is assumed that $M$ takes its extremal value (Eq. (7) in the main text). Since the AdS length $L$ can be scaled away and also the particle mass $m$ enters only through the ratios $E/m$ and $s/m$, we can compute the energy-to-mass ratio of the particle depending on $r_{\text{inj}}$, $s/m$ and $a$. The spin-to-mass ratio $s/m$ is subjected to the point-particle constraint, i.e., $s/m\ll r_{\text{H}}$, while we restrict also to $0\leq a/L<1$.
@@
and use the fact that $E$ and $m$ are constants of motion. Then, when the above quantity is smaller than 111, we have automatically WCCC preservation (according to Fig. 3 in the main text). If, instead, there are some particles that satisfy $\hat{s}/r_{\text{H}}>1$, they live above the dashed line in Fig. 3 and potentially can overspin the BH if they reach the overlapping region when $\hat{s}/r_{\text{H}}>\hat{s}^{\text{AdS}}_{2}/r_{\text{H}}$. A necessary condition for this to happen is that also $\hat{s}/r_{\text{H}}<\hat{s}^{\text{AdS}}_{1}/r_{\text{H}}$ must hold, in order to ensure particle absorption by the BH.
@@
We performed a numerical scan of the parameter space to assess whether such regions, where $\hat{s}/r_{\text{H}}>\hat{s}^{\text{AdS}}_{2}/r_{\text{H}}$, exist. We will consider particles with $s/m\leq 0.2\,r_{\text{H}}$ (so that the point-particle approximation, $s/m\ll r_{\text{H}}$, is satisfied). Fig. 4 reports, for different values of $a/L$, the regions in the parameter space where BH horizon disruption occurs. We have checked that $\hat{s}/r_{\text{H}}<\hat{s}^{\text{AdS}}_{1}/r_{\text{H}}$ is always true in the parameter space considered, so all such particles are absorbed.
@@
Figure 4: Blue shaded regions in the parameter space $[s/(mr_{\text{H}}),\,(r_{\text{inj}}-r_{\text{H}})/r_{\text{H}}]$ (the normalized spin-to-mass ratio and the normalized initial distance of the particle from the horizon, respectively) represent those where the particle has $\hat{s}^{\text{AdS}}_{1}/r_{\text{H}}>\hat{s}/r_{\text{H}}>\hat{s}^{\text{AdS}}_{2}/r_{\text{H}}$ and thus can be absorbed and spin the BH past extremality. Red shaded regions, instead, correspond to regions where the minimum size of the particle (given by $s/m$) is smaller than the distance from the horizon $r_{\text{inj}}-r_{\text{H}}$. As one can see, there is no overlap between the two regions, meaning that particles able to disrupt the horizon actually have a size bigger than the distance from the horizon itself (they are injected already inside the BH). For each plot, we fixed the value of $a/L$ (determining also the value of $r_{\text{H}}$ through Eq. (7) in the main text) and we restricted to particles with $s/m\leq 0.2\,r_{\text{H}}$, starting (with zero velocity) from a radial distance from the horizon equal to $r_{\text{inj}}-r_{\text{H}}\geq 10^{-3}\,r_{\text{H}}$. For $s/m=0$, one always has $\hat{s}/r_{\text{H}}<\hat{s}^{\text{AdS}}_{2}/r_{\text{H}}$, consistent with no horizon disruption.
However, since spinning particles possess a finite size, it is essential to compare this size with their initial distance from the event horizon $r_{\text{inj}}-r_{\text{H}}$. Specifically, the particle’s proper size must be smaller than this initial plunging distance. In Fig. 4, we highlight in red the regions of parameter space where the particle’s size —assumed to be the minimum possible, i.e., $s/m$— is smaller than its initial distance from the horizon. As evident from the figure, these regions do not overlap with those corresponding to particles capable of overspinning the BH (blue shaded region). This leads to the conclusion that particles seemingly able to violate the extremality bound would, in fact, have to be injected behind the horizon and are thus not relevant for the WCCC test.